. Suppose the Y1, Y2, · · · , Yn denote a random sample from a population with Rayleigh distribution (Weibull distribution with parameters 2, θ) with density function f(y|θ) = 2y θ e −y 2/θ, θ > 0, y > 0

Consider the estimators ˆθ1 = Y(1) = min{Y1, Y2, · · · , Yn}, and ˆθ2 = 1 n Xn i=1 Y 2 i .

ii) (10 points) Determine if ˆθ1 and ˆθ2 are unbiased estimators, and in the case that they are not find a multiple of them that make them unbiased.

iii) (5 points) Find the efficiency between the unbiased estimators

iv) (5 points) Are the unbiased estimators consistent?Problem 5. Suppose the Yı, Y2,...,Ydenote a random sample from a population with Rayleigh distribution (Weibull distribution
Answer #1

. Given that Solution is * Suppose the You Y Yo denote a random Sample from a popielation with Reyleigh distoi- bution (Weibu* Fly/B) = 360 et dt. ci-e9%. + The distribution function of yes 70 F(%) - leto | y>0,070. . Let u = y2 =) du = ay dy. u sinccon caso in só e=lo The PDF. of u is . (f(u) = -4%) uxo E (4) flu du. - 99 e/o du. } wat el du 8P8yo ato Lola od 1 Elul. 20:Therefore, day? follows exponential distribution with mean O, consider the eistimators, 8 = Yu = min 2%, Y , . . . , Yn} and:/F(Y) = 1-0 \mus, vertete ang The PDF of 8,743 – dey, (dog) - n beter te como base e - ELM;) fungi, rather this e ampio tuliQue Po Sumtemde CU e Cow] Hence, eev is not unbiased estimator of o, (ii) The efficiency between the unbiased estimator El me20 1 l z. enz/o dz. * MIX) = -79-603-60%) (-). * [44] =% 3 5. #diy % Tyysteryo.uy.7.2010 dz. z² é 20 dzi = 2.0? * v[72) - E[74] - [E(72))? = 20202 * vt. Ex??). Doe The efficiency of a wirrt og is . :Nor como l var Cê, o es (0) colore u-Ji 40 . : moroace 4- J var Cê J. Hence Efficiency of o, wirit is R1 warlo) varrógust H0since, au du zdr. -9%. * F (j) seX dx fly) aine * PDF of 67 = Yong 1. By (Mon) = (er contres, e co - S, () - sob. o og e[ro]Jante te di e netur dy since hy? du nyun su; smero u. arilyn) dyer du: ou sit amet col de Po ju kole4 de . Hence - pence; uwaboot consistorey . Let us consider, EC@ ] ct 2 (132) X no - Elên) = 0. * Vo E ( 6 ) 26 - Et vor (@ny ] -előa) (60))? vor (ona hvar (yo 2) n2 • hvar(yp 2) n2 . or cien) - * lim var lêne ) -0. n > is consistent